### v1

# PACE: Pose Annotations in Cluttered Environments

Abstract
Pose estimation is a crucial task in computer vision, enabling tracking and manipulating objects in images or videos. While several datasets exist for pose estimation, there is a lack of large-scale datasets specifically focusing on cluttered scenes with occlusions. This limitation is a bottleneck in the development and evaluation of pose estimation methods, particularly toward the goal of real-world application in environments where occlusions are common. Addressing this, we introduce PACE (Pose Annotations in Cluttered Environments), a large-scale benchmark designed to advance the development and evaluation of pose estimation methods in cluttered scenarios. PACE encompasses 54,945 frames with 257,673 annotations across 300 videos, covering 576 objects from 44 categories and featuring a mix of rigid and articulated items in cluttered scenes. To annotate the real-world data efficiently, we developed an innovative annotation system utilizing a calibrated 3-camera setup. We test state-of-the-art algorithms in PACE along two tracks: pose estimation, and object pose tracking, revealing the benchmark’s challenges and research opportunities. We plan to release PACE as a public evaluation benchmark, along the annotations tools we developed, to stimulate further advancements in the field. Our code and data is available on https://github.com/qq456cvb/PACE.
1 Introduction
The field of 3D object pose estimation is integral to a myriad of applications, particularly within robotic manipulation. Recent advancements in both instance and category-level pose estimation have been significant, bolstered by deep learning approaches, and perhaps more importantly, data.
PoseCNN [45] advanced pose estimation into the deep learning era, and simultaneously introduced the influential YCB-Video dataset. This benchmark has catalyzed methodological development and offered a consistent evaluation platform. Additionally, the Benchmark for 6D Object Pose Estimation (BOP) challenges have consolidated datasets and refined evaluation metrics for instance-level pose estimation.
In parallel, the NOCS [41] dataset has addressed category-level pose estimation, albeit with a smaller real-world dataset for validation and testing. Despite these strides, the field grapples with a fundamental challenge: existing evaluation datasets are too constrained to thoroughly benchmark the capabilities of pose estimation algorithms. The NOCS REAL275 dataset, for instance, spans only six categories and includes a mere 18 videos. This limitation has led to a performance saturation on current benchmarks, as depicted in Table 1, rendering it ambiguous whether methods are getting tuned to the dataset or improving in general.
| Bottle | Bowl | Camera | Can | Laptop | Mug | |
| SAR-Net [24] | 54.0 | 66.0 | 0.4 | 62.2 | 50.0 | 21.2 |
| HS-Pose [47] | 51.1 | 94.7 | 3.9 | 75.4 | 85.2 | 26.4 |
In this work, we introduce PACE (Common Objects with Pose Annotations), a benchmark for pose estimation, and present a comprehensive study evaluating a wide range of pose estimation and tracking methods. Our contributions are threefold:
The PACE dataset: This dataset includes 576 objects across 44 categories, captured in 300 video clips within diverse scenes. With an average of 183 frames per clip, the dataset encompasses 54,945 frames and 257,673 annotations, providing a large-scale benchmark for pose estimation.
Our evaluation study: To the best of our knowledge, we are the first to analyze and report the performance of state-of-the-art pose estimation methods in a large-scale cluttered setting. These results provide valuable insights that the scalability and generalizability of state-of-the-art methods, making it clear that they are far from reliable in general.
Our annotation pipeline: We will open-source our annotation pipeline, which harnesses a calibrated 3-camera system, enhancing the precision and scalability of annotating poses in real data. This tool significantly mitigates human error and reduces the effort required for annotating 3D poses, providing a solution to one of the major bottlenecks in pose dataset creation.
Overall, our work aims to propel the development of more robust and generalizable pose estimation techniques, thereby facilitating progress towards successful pose estimation in the real world.
| Modality | Cat. | Obj. | Vid. | Img. | Anno. | CAD | Dyn. | Occ. | Marker-free | Artic. | Piled | |
| YCB-Video [45] | RGBD | 1 | 21 | 12 | 20K | 99K | ✓ | ✗ | ✓ | ✓ | ✗ | ✓ |
| LINEMOD-O [3] | RGBD | 1 | 8 | 1 | 1.2K | 9.2K | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ |
| NAVI [16] | RGBD | 1 | 36 | 324 | 10K | 10k | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| NOCS-REAL275 [41] | RGBD | 6 | 42 | 18 | 8K | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | |
| Wild6D [46] | RGBD | 5 | 1722 | 5166 | 1.1M | 1.1M | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ |
| Objectron [1] | RGB | 9 | 17k | 14k | 4M | 4M | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ |
| CO3D [32] | RGB | 50 | 19K | 19K | 1.5M | 1.5M | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| Scan2CAD [2] | RGBD | 9 | 3K | 1506 | - | 14K | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| Pix3D [36] | RGBD | 9 | 395 | - | 10K | 10K | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| HOI-4D [29] | RGBD | 16 | 800 | 4K | 2.4M | - | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
| HANDAL [11] | RGB | 17 | 212 | 2K | 308K | 308K | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| PACE (Ours) | RGBD | 44 | 576 | 300 | 55K | 258K | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
2 Related Works
The field of 3D object pose estimation has seen substantial progress over the past few years. This progress has been facilitated by the introduction of standardized datasets and the development of innovative algorithms.
2.1 Object Pose Datasets
Instance-Level Pose Datasets
YCB-Video dataset [45] is a comprehensive resource for 6D object pose estimation, containing a large number of video frames with accurate pose annotations for 21 objects. LINEMOD-Occluded dataset [3] offers a challenging setting for pose estimation with piled multiple objects in occluded scenes. NAVI dataset [16] presents casually captured images of objects with high-quality 3D scans and precise 2D-3D alignments for advanced 3D reconstruction tasks.
Category-Level Pose Datasets
CO3D [32] offers 1.5 million frames from nearly 19,000 videos across 50 MS-COCO categories for category-specific 3D reconstruction and view synthesis. The Scan2CAD [2] dataset aligns 14225 CAD models from ShapeNet to 1506 ScanNet scans, promoting CAD model alignment in RGB-D scans. Pix3D [36] is a benchmark with image-shape pairs and pixel-level 2D-3D alignment, aiding in shape reconstruction and retrieval. The HOI-4D dataset [29], with 2.4M RGB-D frames over 4000 sequences, enables research in category-level human-object interaction. HANDAL [11] focuses on pose estimation and affordance prediction for robotics-ready manipulable objects. A comparative comparison with other datasets is in Table 2.
2.2 Pose Estimation Methods
Instance-level Pose Estimation
PPF (Point Pair Features) [8] set the pre-deep learning standard for instance-level pose estimation, utilizing local geometric features from point clouds. The deep learning era began with PoseCNN [45], leading to several advanced methods. DeepIM [23] approached pose estimation as an image matching task, iteratively refining estimations. DenseFusion [38] combined global and local features for pose estimation in cluttered scenes, while CosyPose [19] integrated a global refinement strategy in its end-to-end pipeline. SurfEmb [12] leveraged surface embeddings for correspondence matching, and GDRNPP [40] employed geometry-guided regression for enhanced prediction.
Category-level Pose Estimation
Category-level pose estimation extends the challenge to generic object categories. NOCS [41] introduced a unified coordinate space for all objects, predicting object NOCS maps from RGB images. SGPA [4] aims to adapt the structure-guided prior in the pose estimation process, while SAR-Net [24] using shape alignment and symmetric correspondence to estimate a coarse 3D object shape and facilitate object center and size estimation. Recently, HS-Pose [47] proposes a network structure with an HS-layer that extends 3D graph convolution to extract hybrid scope latent features from point clouds for category-level object pose estimation.
2.3 Pose Tracking Methods
Instance-level Pose Tracking Methods like RBOT [37] use RGB data and 3D models to track multiple objects, employing color histograms in their cost function. PoseRBPF [5] separates rotation and translation, using an autoencoder for rotation feature embeddings. ICG [34] iteratively refines pose using geometric cues and is effective for textureless objects, with extensions incorporating visual data [35]. The first deep learning tracker, D6DT [9], and se(3)-TrackNet [43] predict frame-to-frame relative poses, using a render-and-compare strategy.
Category-level Pose Tracking
6-PACK [39] marks the onset of category-level tracking, using DenseFusion [38] features and an attention mechanism for unsupervised keypoint ordering and interframe motion via keypoint matching. CenterPoseTrack [26] projects 2D keypoints from 3D bounding box vertices, achieving RGB-based scale-invariant tracking. BundleTrack [42] generalizes pose tracking without relying on 3D models, instead using video segmentation and pose graph optimization. CAPTRA [44] tracks 9DoF poses for rigid and articulated objects, with subnetworks for rotation regression and normalized coordinate prediction, facilitating analytical 3D size and translation computation.
3 Construction of PACE
A key contribution of this work is the establishment of a scalable and reliable annotation framework, enabling the collection of large-scale and accurate pose annotations. An overview of the pipeline is depicted in Figure 2.
3.1 Acquisition of 3D Common Object Scans
We begin by digitizing an extensive collection of commonplace objects. These items are categorized into 44 distinct classes, as represented in Figure 3.
The Einscan Pro 2X is utilized for rapid scanning of all objects, typically completing within 5 to 10 minutes per object. To expedite this process, we employ a rotatable platform to acquire multiple viewpoints of the objects. After scanning, objects are manually aligned to a standard pose within a uniform coordinate system. We center its axis-aligned bounding box at the origin and align the bounding box orientation along a common axis within the category. We annotate rotational symmetries with the corresponding rotation matrices. High-resolution meshes are then simplified to lower-resolution for smoother annotation workflows.
Articulated Objects
Diverging from many prior pose estimation datasets, our collection encompasses a very wide set of objects, including articulated objects from the AKB48 [27] dataset, namely: scissors, cutters, clips, and boxes. We adopt the alignment methodology from the original AKB48 dataset, without modification. These objects are segmented into multiple parts with hierarchical relationships, presenting a nuanced challenge for pose estimation.
3.2 RGB-D Sequence Acquisition
We designed and implemented a 3-camera system to aid in data acquisition and annotation, comprising three Intel Realsense 415 RGB-D cameras affixed to a metal framework, as illustrated in the bottom of Figure 2. The advantages of this setup include:
Tripling the data yield.
Reducing ambiguity in pose annotation, especially regarding translation along the depth dimension, by using multi-view imagery to enforce consistency across all views.
Enhancing tracking accuracy of static objects with Aruco markers from all three views, making PnP more stable.
Calibration of Multi-Camera Extrinsic Parameters
We calibrate this multi-camera system through a semi-automatic process. Aruco markers initially intended for calibration proved insufficient for high-accuracy rotation estimation. Hence, we resorted to trifocal tensor estimation, i.e. TFT [17]. The process begins with feature extraction and matching, followed by bundle adjustment to refine the positions of the 3D landmarks and camera poses. For reliable feature matching, we employ the SuperPoint [7] descriptor and SuperGlue [33] matcher, using a stringent threshold for matching. We observe that rotational component of the resulting extrinsic parameters is generally precise, but the translation aspect suffers from scale ambiguity inherent in Structure-from-Motion approaches. We correct this by calibrating the scale against markers, applying the following formula to obtain the optimal scale factor:
where is the TFT predicted translation (up to a scale) from camera to camera , is the marker-calibrated translation in real metric scale. We set the extrinsics of the first camera to be the identity matrix.
In cases of repetitive texture patterns, manual intervention is required to establish reliable feature correspondences due to the limitations of SuperPoint+SuperGlue.
3.3 Annotation of Pose Ground-Truths
Previous methodologies have utilized Aruco markers to automate pose estimation through Perspective-n-Points; however, this approach has two drawbacks:
1.
Markers in the scene detract from realism and compromise dataset integrity: training on marker-augmented imagery may result in overfitting to these artificial patterns.
2.
Marker-based annotations are inapplicable to dynamic objects, thus restricting the method’s utility to static scenarios.
Annotation of Static Object Poses
To address the first issue, we employ markers to automate the annotation of static object poses, and then remove the marker appearances from the dataset. We achieve this with a marker inpainting strategy, detailed as follows. Initially at step 1, we place a marker (Marker 1) somewhere within the camera’s field of view. We then record a short video with this marker in view. In step 2, we place a second marker (Marker 2) at a chosen distance from the first, and record another video. After step 2, we remove Marker 1, and begin the actual object capture process (with only Marker 2 present). After this process, we end up with: (1) frames with Marker 1 only, which clearly depict the surface where Marker 2 will later appear; (2) frames capturing both markers, providing helpful calibration cues; (3) frames with Marker 2 only, which represent our main capture. We use the frames from the first two steps to seamlessly inpaint [30] Marker 2’s area in the main capture, as depicted in Figure 4. We leverage Marker 2 for automated pose tracking, and manually correct the tracking every 40 frames in case of drift.
Annotation of Dynamic Object Poses
For dynamic objects, the traditional marker-based tracking approach is insufficient, as the scene markers do not move in tandem with the objects. In such instances, we harness the capabilities of BundleTrack [42], an advanced RGB image-based tracking algorithm. BundleTrack conducts feature correspondence analysis between successive frames to estimate poses, complemented by a bundle adjustment algorithm to optimize keyframes globally and minimize tracking errors. Despite BundleTrack’s proficiency in approximating poses, it is prone to drift, necessitating the manual adjustment of poses every ten frames to ensure precision. This delicate step represents the most labor-intensive aspect of our annotation pipeline, given the current limitations of state-of-the-art tracking techniques in complex scenarios.
Generation of Segmentation Masks
Upon successful pose annotation, we generate occlusion-aware segmentation masks, using z-buffer rendering techniques. In cases where hand interactions are involved, we employ the SAM [18] model to delineate hand masks, and subsequently subtract these from the previously computed object masks to achieve accurate segmentation.
4 Dataset Statistics
In pursuit of diversity, we placed the objects ten disparate environments, each featuring varying levels and configurations of occlusion. We document 10 video sequences per scene, each sequence encompassing 1 to 5 objects. We captured RGB and depth data using an Intel RealSense D415 camera, with a resolution of . The capture process spans distances ranging from 0.5m to 1.5m from the objects.
4.1 Object Distribution and Diversity
The comprehensive distribution of objects is depicted in Figure 3, showcasing a diverse array of both rigid and articulated models. As demonstrated in Figure 5, the dataset encompasses a broad spectrum of object sizes, with the majority measuring approximately 0.2m along the bounding box diagonal. This includes larger items such as storage bins, adding to the heterogeneity of the dataset.
4.2 Variability in Pose, Occlusion, and Environmental Context
The dataset is characterized by a rich variability in object poses, as illustrated in Figure 6, which outlines the statistical distribution of azimuth and elevation angles, indicating comprehensive spatial coverage.
Moreover, we categorized and analyzed the occlusion levels, as shown in Figure 7, delineating instances of severe, moderate, and minor occlusions. Such classification is crucial for assessing the robustness of pose estimation algorithms against varying degrees of visibility.
A visual representation of the ten distinct environmental settings used for data capture is provided in Figure 8, reflecting the contextual diversity of the dataset.
4.3 Dataset Split for Training, Validation, and Testing
To facilitate a comprehensive and equitable evaluation of pose estimation methodologies, we partition the dataset into validation and test subsets following a 20/80 ratio. This separation ensures that a substantial volume of data is available for rigorous testing.
Additionally, to support research necessitating extensive training datasets, we generate a considerable set of synthetic images utilizing a physically based renderer (PBR) [6]. The training set comprises over 52,000 images, each featuring multiple objects, culminating in a vast number of annotations. Qualitative examples of the synthetic images are presented in Figure 9.
5 Evaluation Benchmarks
State-of-the-art pose estimation algorithms typically decompose the task into two distinct stages: detecting or segmenting the object of interest, followed by pose estimation within the predicted bounding box or mask. While the latter has been the primary focus in literature, often leveraging off-the-shelf detectors, a comprehensive evaluation requires an isolated assessment of each stage. Therefore, we examine the pose estimation result under the assumption of perfect detection to allow for equitable comparison across methods. The object detection performance of current state-of-the-art methods can be found in the supplementary and is not the focus of this dataset. We also introduce a benchmark for object pose tracking, premised on the availability of ground-truth data in the initial video frame.
5.1 Pose Estimation Benchmark
This benchmark is divided into instance-level and category-level pose estimation. The former concerns known instances during training, while the latter involves unknown instances similar to those in the training set.
5.1.1 Instance-Level Pose Estimation
This task demands the prediction of rotation and translation for known instances from the training set.
Metrics:
Adhering to the BOP challenge protocol [15], we utilize Average Recall (AR) of Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD) as our metrics. Detailed computation of these metrics is described in [15]. Objects with visibility fraction less than are skipped for evaluation following the BOP [15] convention. Results are averaged across all instances.
Baselines:
We assess four baselines: PPF [8], CosyPose [19], SurfEmb [12], and GDRNPP [40]. While PPF does not require training data, the others are state-of-the-art methods dependent on additional training on the PBR dataset. To focus on pose estimation performance, we assume that ground-truth instance detections are available.
Result Analysis:
The quantitative analysis is presented in Table 3. Notably, while state-of-the-art methods excel on the BOP benchmarks, they do not perform as well as PPF, which relies on local geometric feature matching in point clouds. The relative success of PPF highlights the potential for improvement in state-of-the-art methods, particularly regarding robustness and scalability to handle a large set of instances. Qualitative results can be found in the supplementary.
| Modality | Detection | AR | AR | AR | AR | |
| PPF [8] | D | G.T. | 35.3 | 42.7 | 49.3 | 42.4 |
| CosyPose [19] | RGB | G.T. | 1.4 | 0.3 | 11.5 | 4.4 |
| SurfEmb [12] | RGB | G.T. | 6.2 | 3.0 | 17.8 | 9.0 |
| GDRNPP [40] | RGB-D | G.T. | 3.6 | 2.1 | 15.4 | 7.0 |
| Detection | IoU | IoU | AP | ||||||
| 0:20 | 0:60 | 0:5cm | 0:15cm | 0:20 0:5cm | 0:60 0:15cm | ||||
| NOCS [41] | Mask-RCNN | 0.0/0.0 | 0.0/0.0 | 0.0/0.0 | 0.1/0.0 | 2.8/0.0 | 26.1/0.0 | 0.0/0.0 | 0.1/0.0 |
| HS-Pose [47] | G.T. | 32.7/0.2 | 7.3/0.0 | 5.2/0.0 | 7.0/0.4 | 61.4/41.0 | 86.2/79.5 | 4.0/0.0 | 6.6/0.4 |
| SGPA [4] | G.T. | 1.3/0.0 | 0.0/0.0 | 6.7/1.1 | 13.4/9.0 | 18.5/10.4 | 58.4/49.2 | 3.3/0.3 | 11.7/7.7 |
| DualPoseNet [25] | G.T. | 0.1/0.0 | 0.0/0.0 | 5.1/0.0 | 5.3/0.0 | 15.0/34.6 | 53.6/69.4 | 1.4/0.0 | 4.1/0.0 |
| SAR-Net [24] | G.T. | 25.3/0.5 | 1.1/0.0 | 5.2/0.1 | 7.1/2.6 | 37.6/37.1 | 77.3/77.9 | 2.8/0.1 | 6.3/2.6 |
5.1.2 Category-Level Pose Estimation
Category-level pose estimation tasks involve predicting the 3D bounding box dimensions, rotation, and translation of target instances, given only the category information a priori.
Metrics:
We adopt from NOCS [41], calculating mean Average Precision (AP) across predefined angle and translation thresholds. Specifically, AP@0:20 represents AP averaged from 0 to 20 at 1 intervals; AP@0:5cm is averaged from 0cm to 5cm at 0.25cm intervals; AP@0:20,0:5cm combines these angle and translation thresholds. IoU and IoU denote AP for 3D bounding box matches at IoU thresholds of 25 and 50, respectively. Objects with visibility fraction less than are skipped for evaluation following the BOP [15] convention. Results are separately reported for both rigid and articulated objects.
Baselines:
We evaluate six recent category-level pose estimation methods: NOCS [41], HS-Pose [47], SGPA [4], DualPoseNet [25], SAR-Net [24], and ANCSH [22], the latter specifically designed for articulated objects. For adapting rigid-object-focused methods to articulated objects, each movable part is considered a distinct category. For instance, scissors with two movable parts are treated as two separate categories. Ground-truth detections are presumed for each method, except for NOCS, which outputs both instance masks and pose estimations using a unified network. The compared methods all use the depth as input.
Result Analysis:
Quantitative comparisons are listed in Table 4. On rigid objects, HS-Pose shows excellent performance on IoU and IoU, as well as AP metrics for translation, but falls short on rotation AP metrics compared to SGPA, suggesting a proficiency in size and translation prediction but not rotation. SGPA excels in rotation estimation but demonstrates lower performance on bounding box size metrics. For articulated objects, all methods face challenges due to the movable parts, with ANCSH underperforming on the large-scale real-world dataset, indicative of a significant sim-to-real gap. NOCS shows instability, likely due to its reliance on RGB-based coordinate prediction, which lacks robustness and scalability. Qualitative results are available in supplementary.
5.2 Pose Tracking Benchmark
We categorize state-of-the-art pose tracking methods into model-free, which use a 3D CAD model, and model-based, requiring only the initial pose at the first frame.
Metrics:
For model-free pose tracking, we use four metrics following previous work [39]. 1) , the percentage of predictions with rotation error 5° and translation error 5cm. 2) IoU25, the percentage of intersection over union that is larger than between the two 3D bounding boxes with ground-truth size, transformed by the predicted and ground-truth 6D pose, respectively. 3) , mean value of rotation error in degrees. 4) , mean value of translation error in centimeters. Here the last two metrics are respect to IoU25 since objects with IoU are not counted.
For model-based pose tracking, we report the area under curve (AUC) with respect to ADD, ADD-S [45] and ADD(-S). We set the maximum threshold of AUC to be 0.1m [38]. The ADD metric is first introduced in [14] to calculate average per-point distance between two point clouds, transformed by the predicted pose and the ground-truth, respectively. For symmetric objects like bowls, ADD-S metric is introduced to count for the point correspondence ambiguity. The notation ADD(-S) corresponds to computing ADD for non-symmetric objects and ADD-S for symmetric objects. The objects with visibility fraction less than are skipped for evaluation following the BOP [15] convention. Results are reported by averaging over the rigid/articulated categories, separately.
| Modality | ADD | ADD-S | ADD(-S) | |
| RBOT [37] | RGB | 7.1/0.5 | 10.3/0.8 | 7.4/0.5 |
| ICG [34] | RGB-D | 35.6/10.1 | 48.1/15.2 | 38.1/10.1 |
| Training-Free | Modality | 55cm | IoU25 | R | T | |
| BundleTrack [42] | ✓ | RGB | 6.4/11.2 | 9.1/14.1 | 3.2/5.5 | 2.6/0.8 |
| CAPTRA [44] | ✗ | D | 12.9/4.4 | 47.0/18.5 | 20.2/46.7 | 2.1/1.5 |
| CAPTRA [44] | ✗ | D | 12.9/4.4 | 45.8/20.6 | 19.2/40.9 | 2.2/1.5 |
| 6-PACK [39] | ✗ | RGB-D | 9.2/3.9 | 23.1/16.7 | 17.7/33.6 | 2.1/1.2 |
Baselines:
For methods relying on object mask during tracking, we directly used the ground-truth mask. Since the big gap of performance between rigid and articulated objects, we separately reported their average results.
We regarded each part of the articulated objects as independent for all tracking methods except for CAPTRA [44] which treat all parts of an object as a whole. We trained CAPTRA for each category using our synthetic PBR data.
Result Analysis:
Tables 5 and 6 present the performance metrics of state-of-the-art (SOTA) tracking methods on the PACE dataset with both rigid and articulated results separated by slash. The methods exhibit limited success, with ICG [34] achieving the highest AUC with respect to ADD(-S) at only 38.1% for rigid objects and a mere 10.1% for articulated objects, as shown in Table 5. Table 6 underscores the challenge further, with the best 55cm accuracy below 13%, and IoU25 not exceeding 47%, indicating substantial pose estimation errors for over half of the objects.
The convention followed by 6-PACK [39] and BundleTrack [42] disregards objects with IoU when computing rotational and translational errors (R and T). However, in the context of our dataset where methods generally struggle, this approach could mask true performance levels. This discrepancy is exemplified by BundleTrack [42], which reports lower performance in 55cm and IoU25 metrics yet shows seemingly better results for R and T. Such results indicate that current SOTA methods may require significant improvements to handle the complexity presented by the PACE dataset effectively. Qualitative results can be found in the supplementary.
6 Conclusions and Future Work
This work presented a comprehensive benchmark for 3D object pose estimation and tracking through the introduction of the PACE dataset. Our findings reveal that while there have been significant advancements in pose estimation techniques, there exists a substantial performance gap when these methods are applied to real-world, diverse datasets such as PACE. Particularly, current state-of-the-art methods struggle with articulated objects and exhibit limitations in robustness and scalability. This benchmark serves not only as a testament to the progress achieved but also as a clarion call for the research community to address the complexities of real-world applications.
The results from the PACE dataset underscore a pronounced generalization gap, suggesting that existing models may not sufficiently capture the complexities inherent in diverse real-world scenarios. To bridge this gap, future research could explore the potential of larger, more complex models for pose estimation that can encapsulate a wider variety of object features and environmental contexts.
7 Acknowledgements
This work was supported by the National Key Research and Development Project of China (No. 2021ZD0110700), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), Shanghai Qi Zhi Institute, and SHEITC (2018-RGZN-02046). Yang You is also supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University.
Supplementary Material
A Object Detection Benchmark
Though the primary focus of this dataset is the pose estimation, we are interested in how current state-of-the-art detection models perform on our dataset. In practice, the best method can be served as the default detection method when evaluating the pose estimation result.
A.1 Instance-Level Object Detection
Evaluation Metrics:
Consistent with established literature, we employ Average Precision (AP), Average Precision at IoU thresholds of 50% (AP) and 75% (AP), along with Average Recall (AR) as the metrics for evaluation.
Baseline Methods:
The task is categorized into bounding box (BBox) detection and instance segmentation (Mask). We benchmark the performance against three models: Mask R-CNN [13], YOLO-X [10], and MaskDINO [20]. MaskDINO is notable for its claim of providing precise bounding box and instance segmentation results. Models are trained on a provided PBR dataset, treating each instance as an individual ”category“, culminating in 576 unique categories. Results are averaged across all instances in the test set.
Results and Analysis:
Quantitative outcomes are presented in Table 7. We see that YOLO-X is the best in terms of the bounding box detection while MaskDINO is the best in instance segmentation.
A.2 Category-Level Object Detection
Evaluation Metrics:
The metrics for category-level object detection are identical to those used for instance-level detection.
Baseline Methods:
This task is bifurcated into bounding box (BBox) detection and instance segmentation (Mask). We evaluate the performance of YOLO-X [10], MaskDINO [20], and the zero-shot detector GLIP [21]. GLIP is capable of inferring bounding boxes based on textual descriptions of target categories. For instance segmentation, we compare Mask R-CNN [13], MaskDINO [20], and Grounded SAM [28], which can generate instance masks from text prompts in a zero-shot fashion. Non-zero-shot methods are trained on images from the provided PBR training set.
Results and Analysis:
Quantitative findings are detailed in Table 8. YOLO-X again is the winner is detecting bounding boxes and MaskDINO is the winner in instance segmentation. Though zero-shot methods can achieve some results but they are still far behind the supervised methods, partially due to the text ambiguity in describing the objects.
| Type | ZS | AP | AP | AP | AR | |
| YOLO-X [10] | BBox | ✗ | 50.9 | 66.4 | 57.9 | 64.3 |
| MaskDINO (BBox) | BBox | ✗ | 46.8 | 65.3 | 53.3 | 59.7 |
| GLIP [21] | BBox | ✓ | 22.2 | 30.3 | 26.1 | 59.8 |
| Mask R-CNN [13] | Mask | ✗ | 40.9 | 60.3 | 47.8 | 51.4 |
| MaskDINO (Mask) | Mask | ✗ | 42.9 | 65.1 | 47.4 | 54.6 |
| Grounded-SAM [28] | Mask | ✓ | 8.2 | 11.7 | 9.6 | 15.0 |
B Baseline Adaptations
This section delineates the adaptations made to the baseline methodologies enabling their evaluation on our dataset. Unless otherwise stated, the configurations adhere to the defaults specified in their respective originating papers.
B.1 Handling Object Symmetry
For baselines tasked with rotational regression, we address the ambiguity presented by object symmetry, where multiple rotations correspond to a single input by normalizing these rotations into distinct, unambiguous targets, following the method outlined in [31].
B.2 Instance-Level Pose Estimation Baselines
CosyPose [19]
To maintain comparability, we exclude the random background pasting augmentation from CosyPose. Moreover, to accommodate our hardware constraints, we reduce the image resolution by half.
SurfEmb [12]
With SurfEmb, images are decoded into a shared continuous embedding space alongside 3D point embeddings. Due to the extensive variety of objects in our dataset (576 in total), a separate image decoder for each would exceed our GPU’s memory capacity. Consequently, we implement a unified decoder across all objects.
B.3 Pose Tracking Baselines
CAPTRA [44]
We treat all objects as non-symmetric for training purposes, given that CAPTRA’s loss is designed to handle only asymmetry or continuous symmetry along the y-axis. We also limit our scope to articulated objects comprising two parts, aligning with CAPTRA’s fixed-part count assumption. To manage the higher image resolution and variable object size in our dataset compared to NOCS, we decrease the radius parameter when computing the axis-aligned bounding box aabb and downsample the back-projected point clouds prior to ball filtering.
C Qualitative Comparison on Pose Estimation/Tracking
In this section, we present additional illustrations that further elucidate the performance of various baseline methodologies. Specifically, Figure 10 provides a visualization of the efficacy of instance-level pose estimation methods. Similarly, Figure 11 demonstrates the performance of category-level pose estimation techniques. Additionally, Figure 12 and Figure 13 offer visual representations of the performance metrics for model-based and model-free tracking methods, respectively. These visualizations serve to complement the quantitative analyses provided earlier, offering a more comprehensive understanding of each method’s effectiveness.
D Annotation Software and Pipeline Details
Our annotation interface features a primary display that simultaneously presents views from a triple-camera setup. Adjacently, a sidebar is integrated, showcasing the current annotations within the scene, along with a catalog of potential 3D models from our database. Users can initiate the pose estimation by aligning corresponding points on the 3D models with their 2D image counterparts, utilizing a RANSAC-PnP algorithm. Subsequent refinements to the pose are facilitated through incremental rotations and translations, which can be executed via keys.
For enhanced user experience and efficiency, we have designed a tripartite panel system. This allows users to seamlessly toggle between 2D annotation, 3D annotation, and 2D segmentation workflows. As illustrated in Figure 14, this configuration enables real-time previews of annotations in both 2D and 3D, ensuring an intuitive and dynamic annotation process.
E Snapshots for All the Objects
Figure 15 gives the model snapshots for all our collected objects.